OpenStudy (anonymous):

How do you solve 5(k-2)= 0

OpenStudy (anonymous):

you would go thought the order of operations to get k isolated.

OpenStudy (anonymous):

By multiplying 5 by k then 5 by 2?

OpenStudy (anonymous):

so dividing by 5 on each side will get k-2=0 than just add two to each side and get k=2 :)

OpenStudy (anonymous):

Would this be quadratic equations?

OpenStudy (anonymous):

Well, its on my quadratics homework, so I would think so, yes

OpenStudy (anonymous):

haha, yes, since there is the 5(x-2) its considered a factored equation

OpenStudy (anonymous):

Keep in mind if there isn't a variable on the first constant it wouldn't be considered to be equal to 0.

OpenStudy (anonymous):

so what would the answer be?

OpenStudy (anonymous):

Then as nicole said, replace the positive 2 with the variable. 5(2-2)=0 the above would be correct. And your answer would be k = 2 <-- the 2 would be the answer in this case

OpenStudy (anonymous):

Not sure if I explained so you could understand tho.. im horrible at explaining

OpenStudy (anonymous):

Great. Thanks a bunch to both of you

OpenStudy (anonymous):

need anymore help? :)

OpenStudy (anonymous):

No, I understand how to do it now. Thanks again

OpenStudy (anonymous):

No problem. =)

OpenStudy (anonymous):

-7(2y-1)y=0

OpenStudy (anonymous):

would the answer be for that question y = 1/2, 0

OpenStudy (anonymous):

Yep, plug them in to check if they end up to = to 0 and you'd be correct.

OpenStudy (anonymous):

Lol, both are good.

OpenStudy (anonymous):

Haha thanks :)

OpenStudy (anonymous):

No prob. :)